# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@ 7. Conclusion
Limitations
This survey provides a comprehensive overview of studies on automatic detection of misleading climate-related corporate communications and its related tasks as of early 2025. A number of limitations remain: Given the extensive scope of the field, our review concentrated on NLP-based research, providing only limited discussion of topics like ESG prediction, which are predominantly explored through quantitative methods. Furthermore, our review relied primarily on keyword searches on arXiv, Google Scholar, and ResearchGate, complemented by references from identified papers. In this, we may have overlooked relevant studies that use alternative terminology or are published in less accessible venues. Finally, much of greenwashing happens not on the text level, but on the factual level (with claims that contradict data) or on the multi-modal level (across images or videos). These remain out of scope for our survey.
Acknowledgment
We would like to express our sincere gratitude to Nina Godart and AEF Info for their invaluable contributions and knowledge on the topic. We also extend our thanks to Takaya Sekine CFA for his insightful comments. This work was partially supported by the grant ANR-20-CHIA-0012-01 (“NoRDF”).
Appendix
@@ A.1. Data sources
10-K reports (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023)
Manually collected reports (Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022; Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023)
Refinitiv (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022)
Reuters (Li et al., 2020; Clarkson et al., 2020)
ESG score private providers (Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b)
CDP (Spokoyny et al., 2023)
Open for Good Data (Vinella et al., 2023)
ClimateWatch (Kraus et al., 2023)
Twitter/Reddit (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024)
LobbyMap (Morio and Manning, 2023)
Net-zero tracker (Schimanski et al., 2023a)
Fact-checking websites (Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021)
Table 10. Data Sources
@@
Many studies rely on 10-K reports (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023) as they are easily accessible (EDGAR), uniformly structured, and mandatory for companies listed on US financial markets. Other reports, such as sustainable reports, are not centralized and mandatory, therefore they need to be collected from other sources. They are often collected manually from the company’s website (Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022), or from other websites (Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023) such as from the GRI, SASB, TCFD, ResponsibilityReports, AnnualReports, Korean Standards Association. Polignano et al. (2022) relied on Sustainability Disclosure Database from GRI. To the best of our knowledge, it is not available any longer.
The reports are openly available, however they are most of the time in a PDF format, and must be parsed to produce machine-readable versions. While Khan (2024) proposed a machine-readable dataset for the 10-K reports, it is often not the case for other types of reports. (Bronzini et al., 2023; Polignano et al., 2022; Kang and Kim, 2022) provide a sample of processed reports.
@@
Machine-readable reports, as well as earning call transcripts are often provided through a private data provider such as Refinitiv (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022) or Reuters (Li et al., 2020; Clarkson et al., 2020). Similarly, financial data (e.g. ESG score) is often collected through data providers such as RavenPack or S&P Global Trucost or Wharton’s research platform WRDS, Reuters, Bloomberg, Refinitiv, RobecoSAM and Compustat (Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b)
@@
Carbon emissions of companies can be collected through the private data providers, or through CDP (Spokoyny et al., 2023), Open for Good Data (Vinella et al., 2023) or ClimateWatch website (Kraus et al., 2023).
@@
As companies communicate with the public mainly through marketing and social media, other data sources include Twitter and Reddit (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024).
@@
Finally, projects such as LobbyMap or Net Zero Tracker, that track corporate policy and engagements, can help build annotated datasets (Morio and Manning, 2023; Schimanski et al., 2023a). Similarly, relying on a fact-checking website such as https://climatefeedback.org/ or www.skepticalscience.com is an efficient way of building a dataset (Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021)
Appendix B Detailed model list
Table 11. Solutions proposed in each paper
Paper Dataset Baselines BERT Sent. Sim. Other
Pretraining Models on Climate-Related Text
Webersinke et al. (2022) Domain-specific corpus - Pre-Training ClimateBERT
Vaghefi et al. (2022) Deep Climate Change - Pre-Training ClimateGPT-2
Mullappilly et al. (2023) Deep Climate Change - Pre-Training stsb-xlm-r-multilingual Arabic Mini-ClimateGPT, Vicuna, Alpaca, Dolly V2, Alpaca ar
Schimanski et al. (2023b) Domain-specific corpus - Pre-Training EnvironmentalBERT
Yu et al. (2024) climateBUG-data - Pre-Training climateBUG-LM
Thulke et al. (2024) Domain-specific corpus - Pre-Training ClimateGPT-7/13/70B
Thulke et al. (2024) ClimaBENCH - Evaluation Stability-3B, Pythia-6.9B, Falcon-7B, Mistral-7B, Llama-2-7B, Jais-13B, Jais-13B-Chat, Llama-2-7/13/70B, Llama-2-Chat-7/13/70B, ClimateGPT-7/13/70B
Climate-related topic detection
Doran and Quinn (2009) 10-K filings - climate Keywords
Engle et al. (2020) WSJ - climate TF-IDF cosine similarity
Leippold and Varini (2020) climatext NB, Keywords BERT
Garrido-Merchán (Garrido-Merchán et al., 2023) climatext ClimateBERT, BERT
Spokoyny et al. (2023) climatext Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Webersinke et al. (2022) climate detection distilRoBERTa, climateBERT
Bingler et al. (2023) climate detection NB, SVM + BoW, SVM + ELMo climateBERT
Brié et al. (2022) climate detection Random Forest climateBERT, FinBERT
Yu et al. (2024) climateBUG data BERT, FinBERT, ClimateBERT, climateBUG-LM
Lin et al. (2023b) sustainable signals review
Climate-related subtopic detection
Kang and Kim (2022) Sustainable Reports - topics sentence-BERT
Mishra and Mittal (2021) sciDCC
Spokoyny et al. (2023) sciDCC Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Vaid et al. (2022) climateEng BERT, RoBERTa, DistilBERT FastText
Spokoyny et al. (2023) climateEng Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Jain et al. (2023) ledger - topics TF-IDF BERT, ClimateBERT and RoBERTa. sentence-BERT Word2Vec-based Random Forest
Schimanski et al. (2024a) ESGBERT-nature Keywords ClimateBERT, RoBERTa, DistilRoBERTa, EnvironmentalBERT
Spokoyny et al. (2023) ClimaTOPIC Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Detecting climate-related financial disclosure
Ding et al. (2023) Reports - Disclosure TF-IDF cosine similarity
Auzepy et al. (2023) Reports - TCFD BART+MNLI
Brié et al. (2022) Sustainable Reports structural topic model (STM)
Bingler et al. (2022) climate tcfd rec. TF-IDF + random forest RoBERTa, ClimateBERT (RoBERTa+LogReg) Sentence Enc.
Sampson et al. (2022) Reports - TCFD TF-IDF ClimateBERT, distilRoBERTa kNN (sentence-transformers embeddings), stacked models
In-depth Disclosure: Climate Risk Classification
Li et al. (2020) Reports - Risk dictionaries (risk)
Chou et al. (2023) 10-K filings - Risk 3 dictionaries (risk: Climate, Impact, Litigious)
Kheradmand et al. (2021) Annual/Sustainability Reports - Risk dictionaries (risk)
Sautner et al. (2023) earning calls - Risk dictionaries (risk/opportunity)
Polignano et al. (2022) Sustainability Reports - Sentiment TextBlob, Sentita
Kang and Kim (2022) Sustainability Reports - Sentiment DistilBERT (sentiment analysis)
Kölbel et al. (2021) 10-K filings - Risk BoW, TF-IDF BERT
Friederich et al. (2021) Reports - Risk SVM distilBERT, RoBERTa
Bingler et al. (2023) climate sentiment NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Webersinke et al. (2022) climate sentiment distilRoBERTa, climateBERT
Xiang and Fujii (2023) Web of Science, Scopus - Sentiment BERT, Domain-specific BERT, TinyBERT, DistilBERT, ClimateBERT CCLA+Max-Pooling, Bi-LSTM-Attention, POS-Bi-LSTM-Attention, DARE
Detecting environmental, social, and governance disclosure
Rouen et al. (2023) Sustainability Reports - ESG industry-topic dictionaries
Mehra et al. (2022) Sustainability Reports - ESG sentence BERT, Universal Sentence Encoder, and Deep Averaging Network
Huang et al. (2020) Sustainability Report, 10-K filings - ESG NB, SVM, RF FinBERT, BERT CNN, LSTM
Schimanski et al. (2023b) ESGBERT E,S,G BERT, RoBERTa, DistilRoBERTa, E/S/G-RoBERTa, E/S/G-DistilRoBERTa
Lee and Kim (2023) Korean Reports - ESG Multilingual BERT, KoBERT, DistilKoBERT, KB-ALBERT, KLUE-RoBERTa
Bronzini et al. (2023) Sustainability Reports - ESG INSTRUCTOR-xl
Green claim detection
Woloszyn et al. (2021) green claims RoBERTa, BERTweet Flair
Stammbach et al. (2023) environmental claims Random, Majority, TF-IDF + SVM, Character n-gram SVM DistilBERT, ClimateBERT, RoBERTa ClaimBuster RoBERTa, Pledge Detection RoBERTa
Lin et al. (2023b) Product Description - Claim environmental claim detection (finetuned distilBERT)
Green claim characteristics
Bingler et al. (2023) climate-commitments-actions NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Schimanski et al. (2023a) netzero/reduction ClimateBERT, DistilRoBERTa, RoBERTa GPT-3.5-turbo
Woloszyn et al. (2021) green claims 3 RoBERTa, BERTweet Flair
Bingler et al. (2023) climate specificity NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Green stance detection
Morio (Morio and Manning, 2023) LobbyMap Majority, TF-IDF BERT, RoBERTa, ClimateBERT, Longformer
Luo et al. (2020) DeSMOG
Vaid et al. (2022) ClimateStance BERT, RoBERTa, DistilBERT FastText
Lai and Chen (2023) Reports - Stance BERT, RoBERTa, ClimateBERT
Spokoyny et al. (2023) ClimateStance Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Spokoyny et al. (2023) DeSMOG Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Leippold (Leippold and Diggelmann, 2020) ClimateFEVER ALBERT(FEVER)
Wang et al. (2021) ClimateFEVER RoBERTa (ClimateFEVER) RoBERTa (FEVER), RoBERTa (ClimateFEVER) + UDA
Spokoyny et al. (2023) ClimateFEVER Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Vaghefi et al. (2022) ClimateFEVER GPT-2, climateGPT-2
Xiang and Fujii (2023) ClimateFEVER BERT, RoBERTa, DistilRoBERTa, ClimateBERT, POS-Bi-LSTM-Attention
Webersinke et al. (2022) ClimateFEVER distilRoBERTa, climateBERT
Vaid et al. (2022) ClimateFEVER BERT, RoBERTa, DistilBERT FastText
Question Answering
Luccioni et al. (2020) Reports - QA RoBERTa
Spokoyny et al. (2023) ClimaQA BM25 MiniLM
Spokoyny et al. (2023) ClimaINS Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Rony et al. (2022) CCMRC - QA pairs ALBERT (trained on SQuAD then finetuned) Sentence-BERT
Schimanski et al. (2023a) Net-Zero Tracker - Target RoBERTa Quad2
Ni et al. (2023) Reports - RAG ChatGPT
Vaghefi et al. (2023) Reports - RAG - Human Eval GPT-4
Mullappilly et al. (2023) Clima500-Instruct - QA instructions Vicuna (Mini-ClimateGPT)
Thulke et al. (2024) Instruction following corpus - Human Eval ClimateGPT-70B, ClimateGPT-7B, ClimateGPT-FSC-7B
Thulke et al. (2024) ClimRetrieve ColBERTv2, DRAGON, GTE-base, text-embedding-ada-002, text-embedding-3-small, text-embedding-3-large BM25
Deceptive Techniques
Oppong (Oppong-Tawiah and Webster, 2023) Tweets - Fake News linguistic indicators
Coan et al. (2021) contrarian claims Logistic, SVM RoBERTa ULMFiT (over sample, weighted, focal Loss, logisitic, SVM), RoBERTa + Logistic
Jin et al. (2022) logicClimate ELECTRA, BERT lectra-StructureAware
Bhatia et al. (2021) climate change scepticism - Neutralization SVM BERT MTEXT
Environmental performance prediction
Jain et al. (2023) Ledger - Scope 3 Emissions Alpaca, WizardLM, Flan-T5, ChatGPT
Bronzini et al. (2023) Sustainable Reports - ESG OLS
Mehra et al. (2022) Reports 10-Qs - ESG Majority BERT, ESGBERT
Clarkson et al. (2020) CSR Reports - CSR performer linguistic features + random forest, XGBoost, SVM
Lin et al. (2023b) Product Description / Review - Finch Score Lasso, Gradient Boosting NOCATE + transformer, SUSTAINABLESIGNALS + transformer
Sharma et al. (2024) Multimodel Tweet - ESG CLIP+MLP, Sentence-BERT+MLP
@@ Appendix C Reported Performances
C.1. Climate-related Topic Classification
Source Model F1
Webersinke et al. (2022) distilRoBERTa 98.6% ±$\pm 1$
Webersinke et al. (2022) ClimateBERT 99.1% ±$\pm 0.8$
Bingler et al. (2023) Naive Bayes 87%
Bingler et al. (2023) SVM + BoW 87%
Bingler et al. (2023) SVM + ELMo 89%
Bingler et al. (2023) ClimateBERT 97%
Brié et al. (2022) ClimateBERT 98.59%
Brié et al. (2022) FinBERT 96.67%
Brié et al. (2022) Random Forest 91.65%
(a) F1-scores (weighted average) on ClimateBERT’s climate detection dataset (Bingler et al., 2023). For Webersinke et al. (2022) scores are displayed as mean ±$\pm std$. Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
Source Model F1
Yu et al. (2024) BERT 90.81%
Yu et al. (2024) FinBERT 90.82%
Yu et al. (2024) ClimateBERT 91.07%
Yu et al. (2024) climateBUG-LM 91.36%
(b) F1-scores (macro average) on ClimateBUG-data (Yu et al., 2024).
@@ C.1. Climate-related Topic Classification
F1 (bin.) F1 (bin.) F1 (bin.)
Leippold and Varini (2020) best BERT 80% ±$\pm 6$ 95% $\pm$ 2 83% $\pm$ 1
Leippold and Varini (2020) best NB 60% ±$\pm 7$ 90% ±$\pm 3$ 72% ±$\pm 2$
Leippold and Varini (2020) best Keywords 67% ±$\pm 7$ 88% ±$\pm 3$ 70% ±$\pm 1$
Leippold and Varini (2020) BERT-AL-wiki 69% ±$\pm 6$ 71% ±$\pm 4$ 81% ±$\pm 1$
Garrido-Merchán (Garrido-Merchán et al., 2023) BERT-AL-wiki - 91% ±$\pm 0.6$
Garrido-Merchán (Garrido-Merchán et al., 2023) climateBERT-AL-wiki - 93% ±$\pm 0.7$
Combined - F1 (macro)
Spokoyny et al. (2023) Majority 42.08%
Spokoyny et al. (2023) Random 46.86%
Spokoyny et al. (2023) SVM 83.39%
Spokoyny et al. (2023) BERT 87.04%
Spokoyny et al. (2023) RoBERTa 85.97%
Spokoyny et al. (2023) DistilRoBERTa 86.06%
Spokoyny et al. (2023) Longformer 87.80%
Spokoyny et al. (2023) SciBERT 83.29%
Spokoyny et al. (2023) ClimateBERT 85.14%
(c) F1-scores on ClimaText (Leippold and Varini, 2020). For Leippold and Varini (2020) scores are displayed as mean ±$\pm std$.
Table 12. Reported performances for the datasets on climate-related topic detection (which include climate, sustainability and environment).
C.2. Thematic Analysis
Source Model Prec.
Bingler et al. (2022) TF-IDF+random forest 24%
Bingler et al. (2022) Sentence Enc. 23%
Bingler et al. (2022) RoBERTa Para. 22%
Bingler et al. (2022) RoBERTa Sent. 75%
Bingler et al. (2022) ClimateBERT* 84%
Bingler et al. (2022) ClimateBERT* blahblF1: 84%
(a) Average Precisions and F1-score (macro average) on TCFD recommendations classification (Bingler et al., 2022). (*They refer to a finetuned RoBERTa with a Logistic Regression as ClimateBERT)
@@ C.2. Thematic Analysis
Sampson et al. (2022) TF-IDF (1) 0.867
Sampson et al. (2022) kNN (2) 0.847
Sampson et al. (2022) ClimateBERT (3) 0.852
Sampson et al. (2022) distilRoBERTa (4) 0.819
Sampson et al. (2022) (1) + (2) 0.884
Sampson et al. (2022) (1) + (3) 0.865
Sampson et al. (2022) (1) + (4) 0.857

### v2

@@ 7 Conclusion
[…]
Acknowledgements
We are grateful to Takaya Sekine for his insightful feedback on the manuscript. We further acknowledge Amundi’s climate analysts, notably Aaron Macdougall, and the journalists at AEF Info, in particular Nina Godart, for their assistance in researching the topic and refining definitions.
Declarations
Funding
This work is funded by Amundi Technology, Inria, Télécom Paris, and ANRT through a CIFRE partnership program.
Conflicts of interest/Competing interests
Aside from the authors’ employers—Amundi Technology, Inria, and Télécom Paris—there are no conflicts of interest related to the present work.
Ethics approval and consent to participate
Not applicable.
Consent for publication
Not applicable.
Data availability
Not applicable.
Materials availability
Not applicable.
Code availability
Not applicable.
Author contribution
This work was primarily conducted by Tom Calamai, particularly the research efforts (collection and analysis of papers). All authors contributed to the writing of the manuscript.
Preprint
An earlier version of this work is available as a preprint on arXiv [calamai2025corporategreenwashingdetectiontext].
@@ Appendix B Data sources
10-K reports [doran_risk_disclosure, kolbel_ask_2021, chou_ESG, huangFinBERTLargeLanguage2020, krausEnhancingLargeLanguage2023]
Manually collected reports [kheradmand2021a, rouenEvolutionESGReports2023, lai_using_2023, hyewon_kang_analyzing_2022, auzepy_evaluating_2023, bjarne_brie_mandatory_2022, nicolas_webersinke_climatebert_2021, luccioni_analyzing_2020, kheradmand2021a, bronzini_glitter_2023, LEE2023119726, marco_polignano_nlp_2022, krausEnhancingLargeLanguage2023]
Refinitiv [SAUTNER_cliamte_change_exp, bingler2023cheaptalkspecificitysentiment, stammbach_environmental_2023, Friederich_climate_risk_disclosure, bjarne_brie_mandatory_2022, LEE_greenwashing, nicolas_webersinke_climatebert_2021]
Reuters [liCorporateClimateRisk2020, clarkson_nlp_us_csr]
ESG score private providers [liCorporateClimateRisk2020, SAUTNER_cliamte_change_exp, Mehra_2022, Greenscreen, clarkson_nlp_us_csr, schimanski_bridging_2023]
CDP [spokoyny2023answering]
Open for Good Data [avalon_vinella_leveraging_2023]
ClimateWatch [krausEnhancingLargeLanguage2023]
Twitter/Reddit [vinicius_woloszyn_towards_2021, luo_detecting_2020, vaid-etal-2022-towards, coanComputerassistedClassificationContrarian2021, divinus_oppong-tawiah_corporate_2023, Greenscreen]
LobbyMap [morio2023an]
Net-zero tracker [tobias_schimanski_climatebert-netzero_2023]
Fact-checking websites [jin-etal-2022-logical, varini_climatext_2020, coanComputerassistedClassificationContrarian2021]
Table 12: Data Sources
@@
Many studies rely on 10-K reports [doran_risk_disclosure, kolbel_ask_2021, chou_ESG, huangFinBERTLargeLanguage2020, krausEnhancingLargeLanguage2023] as they are easily accessible (EDGAR), uniformly structured, and mandatory for companies listed on US financial markets. Other reports, such as sustainable reports, are not centralized and mandatory, therefore they need to be collected from other sources. They are often collected manually from the company’s website [kheradmand2021a, rouenEvolutionESGReports2023, lai_using_2023, hyewon_kang_analyzing_2022, auzepy_evaluating_2023, bjarne_brie_mandatory_2022, nicolas_webersinke_climatebert_2021], or from other websites [luccioni_analyzing_2020, kheradmand2021a, bronzini_glitter_2023, LEE2023119726, marco_polignano_nlp_2022, krausEnhancingLargeLanguage2023] such as from the GRI, SASB, TCFD, ResponsibilityReports, AnnualReports, Korean Standards Association. marco_polignano_nlp_2022 relied on Sustainability Disclosure Database from GRI. To the best of our knowledge, it is no longer available.
The reports are openly available; however, they are most of the time in a PDF format, and must be parsed to produce machine-readable versions. While financial-reports-sec proposed a machine-readable dataset for the 10-K reports, it is often not the case for other types of reports. [bronzini_glitter_2023, marco_polignano_nlp_2022, hyewon_kang_analyzing_2022] provide a sample of processed reports.
@@
Machine-readable reports, as well as earning call transcripts are often provided through a private data provider such as Refinitiv [SAUTNER_cliamte_change_exp, bingler2023cheaptalkspecificitysentiment, stammbach_environmental_2023, Friederich_climate_risk_disclosure, bjarne_brie_mandatory_2022, LEE_greenwashing, nicolas_webersinke_climatebert_2021] or Reuters [liCorporateClimateRisk2020, clarkson_nlp_us_csr]. Similarly, financial data (e.g. ESG score) is often collected through data providers such as RavenPack or S&P Global Trucost or Wharton’s research platform WRDS, Reuters, Bloomberg, Refinitiv, RobecoSAM and Compustat [liCorporateClimateRisk2020, SAUTNER_cliamte_change_exp, Mehra_2022, Greenscreen, clarkson_nlp_us_csr, schimanski_bridging_2023]
@@
Carbon emissions of companies can be collected through the private data providers, or through CDP [spokoyny2023answering], Open for Good Data [avalon_vinella_leveraging_2023] or ClimateWatch website [krausEnhancingLargeLanguage2023].
@@
As companies communicate with the public primarily through marketing and social media, other data sources include Twitter and Reddit [vinicius_woloszyn_towards_2021, luo_detecting_2020, vaid-etal-2022-towards, coanComputerassistedClassificationContrarian2021, divinus_oppong-tawiah_corporate_2023, Greenscreen].
@@
Finally, projects such as LobbyMap or Net Zero Tracker, which track corporate policy and engagements, can help build annotated datasets [morio2023an, tobias_schimanski_climatebert-netzero_2023]. Similarly, relying on a fact-checking website such as https://climatefeedback.org/ or www.skepticalscience.com is an efficient way of building a dataset [jin-etal-2022-logical, varini_climatext_2020, coanComputerassistedClassificationContrarian2021]
@@ Appendix C Reported Performances
Climate-related Topic Classification
@@
binary F1 F1 F1
varini_climatext_2020 best BERT 80% $\pm$ 6 95 % $\pm$ 2 83% $\pm$ 1
varini_climatext_2020 best NB 60% $\pm$ 7 90 % $\pm$ 3 72 % $\pm$ 2
varini_climatext_2020 best Keywords 67% $\pm$ 7 88% $\pm$ 3 70% $\pm$ 1
varini_climatext_2020 BERT-AL-wiki 69% $\pm$ 6 71% $\pm$ 4 81 % $\pm$ 1
garridomerchán2023finetuning BERT-AL-wiki - 91% $\pm$ 0.6 -
garridomerchán2023finetuning climateBERT-AL-wiki - 93% $\pm$ 0.7 -
Combined
F1 (macro)
spokoyny2023answering Majority 42.08%
spokoyny2023answering Random 46.86%
spokoyny2023answering SVM 83.39%
spokoyny2023answering BERT 87.04%
spokoyny2023answering RoBERTa 85.97%
spokoyny2023answering DistilRoBERTa 86.06%
spokoyny2023answering Longformer 87.80%
spokoyny2023answering SciBERT 83.29%
spokoyny2023answering ClimateBERT 85.14%
(a) Reported performance on ClimaText [varini_climatext_2020] (annotated datasets).
Source Model F1 (macro)
yu_climatebug_2024 BERT 90.81%
yu_climatebug_2024 FinBERT 90.82%
yu_climatebug_2024 ClimateBERT 91.07%
yu_climatebug_2024 climateBUG-LM 91.36%
(b) Reported Performances on ClimateBUG-data yu_climatebug_2024.
Source Model F1 (weighted average)
nicolas_webersinke_climatebert_2021 distilRoBERTa 98.6% $\pm$ 1
nicolas_webersinke_climatebert_2021 ClimateBERT 99.1% $\pm$ 0.8
bingler2023cheaptalkspecificitysentiment Naive Bayes 87%
bingler2023cheaptalkspecificitysentiment SVM + BoW 87%
bingler2023cheaptalkspecificitysentiment SVM + ELMo 89%
bingler2023cheaptalkspecificitysentiment ClimateBERT 97%
bjarne_brie_mandatory_2022 ClimateBERT 98.59%
bjarne_brie_mandatory_2022 FinBERT 96.67%
bjarne_brie_mandatory_2022 Random Forest 91.65%
(c) Performances on ClimateBERT’s climate detection dataset bingler2023cheaptalkspecificitysentiment. bingler2023cheaptalkspecificitysentiment also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
Table 13: Reported performance for the datasets on climate-related topic detection (which include climate, sustainability and environment)
Thematic Analysis
Source Model Average P F1 (macro)
bingler_cheap_2021 TF-IDF+random forest 24 % -
bingler_cheap_2021 Sentence Enc. 23% -
bingler_cheap_2021 RoBERTa Para. 22% -
bingler_cheap_2021 RoBERTa Sent. 75% -
bingler_cheap_2021 ClimateBERT (RoBERTa+LogReg) 84% 84%
(a) Performance on TCFD recommendations classification [bingler_cheap_2021] with 5 labels
@@
sampson_tcfd-nlp_nodate TF-IDF (1) 0.867
sampson_tcfd-nlp_nodate kNN (2) 0.847
sampson_tcfd-nlp_nodate FT ClimateBERT (3) 0.852
sampson_tcfd-nlp_nodate FT distilRoBERTa (4) 0.819
sampson_tcfd-nlp_nodate (1) + (2) 0.884
sampson_tcfd-nlp_nodate (1) + (3) 0.865
sampson_tcfd-nlp_nodate (1) + (4) 0.857